XYZ company will pay dividend next year of $4.20 per share. The growth rate is constant at 6% and the required return is 10%. What is the price of abc stock at year 5?
Answer #1

Total = Value of stock today = 111.30

Year

Growth

Dividend Cash flow

Price of stock*

Net cash inflow

Discounting = Df

Present value

n

g

Dn = D0*(1+g)^n

Price of stock* = Dividend at end of period x (1+Constant growth)/(Required rate -Constant growth)

NF = Dn + P

Df = 1/(1+10%)^Year

=NF*Df

1

6.00%

                                                                          4.45200000

                  4.45200

                0.909091

4.04727

2

6.00%

                                                                          4.71912000

                  4.71912

                0.826446

3.90010

3

6.00%

                                                                          5.00226720

                  5.00227

                0.751315

3.75828

4

6.00%

                                                                          5.30240323

                  5.30240

                0.683013

3.62161

5

6.00%

                                                                          5.62054743

              148.94451

              154.56505

                0.620921

95.97274

Total = Value of stock today =

111.30

Price of stock* = Dividend at end of period x (1+Constant growth)/(Required rate -Constant growth)

= 5.62054743 x (1+6%)/(10%-6%)

= 148.94451